OpenStudy (anonymous):

if the probability of surviving a head on car accident at 55 mph is 0.059 then what is the probability of not surviving

OpenStudy (anonymous):

\[1-0.059\]

jimthompson5910 (jim_thompson5910):

P(surviving) = probability of surviving P(not surviving) = probability of not surviving P(surviving) + P(not surviving) = 1 P(not surviving) = 1 - P(surviving) P(not surviving) = 1 - 0.059 So that's how satellite73 got " 1 - 0.059 "